### v1

# Amotz Bar-Noy

Degree Realization by Bipartite Multigraphsthanks: A preliminary version of this paper appeared in SIROCCO 2023 [4]. A preliminary version of one of the results of this paper appeared in SWAT 2022 [2]. This work was supported by US-Israel BSF grant 2022205.
Amotz Bar-Noy
amotz@sci.brooklyn.cuny.edu
The City University of New York (CUNY), NY 10016, USA.
   Toni Böhnlein33footnotemark: 3
toni.bohnlein@weizmann.ac.il
   David Peleg
david.peleg@weizmann.ac.il
The Weizmann Institute of Science, Rehovot 76100, Israel.
   Dror Rawitz
dror.rawitz@biu.ac.il
Bar Ilan University, Ramat-Gan 52900, Israel. Part of the research was done while this author was on sabbatical at the department of computer science of Reykjavik University.
The problem of realizing a given degree sequence by a multigraph can be thought of as a relaxation of the classical degree realization problem (where the realizing graph is simple). This paper concerns the case where the realizing multigraph is required to be bipartite.
The problem of characterizing degree sequences that can be realized by a bipartite (simple) graph has two variants. In the simpler one, termed BDRP, the partition of the degree sequence into two sides is given as part of the input. A complete characterization for realizability in this variant was given by Gale and Ryser over sixty years ago. However, the variant where the partition is not given, termed BDR, is still open.
For bipartite multigraph realizations, there are also two variants. For BDRP, where the partition is given as part of the input, a complete characterization was known for determining whether there is a multigraph realization whose underlying graph is bipartite, such that the maximum number of copies of an edge is at most $r$. We present a complete characterization for determining if there is a bipartite multigraph realization such that the total number of excess edges is at most $t$. We show that optimizing these two measures may lead to different realizations, and that optimizing by one measure may increase the other substantially. As for the variant BDR, where the partition is not given, we show that determining whether a given (single) sequence admits a bipartite multigraph realization is NP-hard. On the positive side, we provide an algorithm that computes optimal realizations for the case where the number of balanced partitions is polynomial, and present sufficient conditions for the existence of bipartite multigraph realizations that depend only on the largest degree of the sequence.
Keywords: Degree Sequences, Graph Realization, Bipartite Graphs, Graphic Sequences, Bigraphic Sequences, Multigraph Realization.
1 Introduction
1.1 Background and Motivation
Degree realization.
This paper concerns a classical network design problem known as the Graphic Degree Realization problem (GDR). The number of neighbors or connections of a vertex in a graph is called its degree, and it provides information on its centrality and importance. For the entire graph, the sequence of vertex-degrees is a significant characteristic which has been studied for over sixty years. The graphic degree realization problem asks if a given sequence of positive integers $d=(d_{1},...,d_{n})$ is graphic, i.e., if it is the sequence of vertex-degrees of some graph. Erdös and Gallai [8] gave a characterization for graphic sequences, though not a method for finding a realizing graph. Havel and Hakimi [12, 11] proposed an algorithm that either generates a realizing graph or proves that the sequence is not graphic.
Relaxed degree realization by multigraphs.
An interesting direction in the study of realization problems involves relaxed (or approximate) realizations (cf. [1]). Such realizations are well-motivated by applications in two wider contexts. In scientific contexts, a given sequence may represent (noisy) data resulting from an experiment, and the goal is to find a model that fits the data. In such situations, it may happen that no graph fits the input degree sequence exactly, and consequently it may be necessary to search for the graph “closest” to the given sequence. In an engineering context, a given degree sequence constitutes constraints for the design of a network. It might happen that satisfying all of the desired constraints simultaneously is not feasible, or causes other issues, e.g., unreasonably increasing the costs. In such cases, relaxed solutions bypassing the problem may be relevant.
In the current paper we focus on a specific type of relaxed realizations where the graph is allowed to have parallel edges, namely, the realization may be a multigraph. It is easy to verify that if (multiple) self-loops are allowed, then every sequence $d=(d_{1},\ldots,d_{n})$ whose sum $\sum_{i}d_{i}$ is even has a realization by a multigraph. Hence, we focus on the case where self-loops are not allowed.
The problem of degree realization by multigraphs has been studied in the past as well. Owens and Trent [15] gave a condition for the existence of a multigraph realization. Will and Hulett [19] studied the problem of finding a multigraph realization of a given sequence such that the underlying graph of the realization contains as few edges as possible. They proved that such a realization is composed of components, each of which is either a tree or a tree with a single odd cycle. Hulett, Will, and Woeginger [13] showed that this problem is strongly NP-hard.
Degree realization by bipartite graphs.
The Bigraphic Degree Realization problem (BDR) is a natural variant of the graphic degree realization problem, where the realizing graph is required to be bipartite. The problem has a sub-variant, denoted BDRP, in which two sequences are given as input, representing the vertex-degree sequences of the two sides of a bipartite realizing graph. (In contrast, in the general problem, a single sequence is given as input, and the goal is to find a realizing bipartite graph based on some partition of the given sequence.) The BDRP problem was solved by Gale and Ryser [9, 17] even before Erdos and Gallai’s characterization of graphic sequences. However, the general problem – mentioned as an open problem over forty years ago [16] – remains unsolved today.
A sequence of integers $d=(d_{1},...,d_{n})$ can only be bigraphic, i.e., the vertex-degree sequence of a bipartite graph, if it can be partitioned into two sub-sequences or blocks of equal total sum. The later problem is known as the partition problem and it is solvable in polynomial time assuming that $d_{1}<n$ (which is a necessary condition for $d$ to be bigraphic). Yet, BDP bears two obstacles. First, a sequence may have several partitions of which some are bigraphic and others are not. Second, the number of partitions may be exponentially large in $n$. Recent attempts on the BDR problem (see [2, 3]) try to identify a small set of partitions, which are suitable to decide BDR for the whole sequence. Each partition in the small set is tested using the Gale-Ryser characterization. In case all of them fail the test, it is conjectured that no partition of the sequence is bigraphic. The conjecture was shown to be true in case there exists a special partition that (perfectly) splits the degrees into small and large ones.
Paralleling the above discussion concerning relaxed degree realizations by general multigraphs, one may look for relaxed degree realizations by bipartite multigraphs. This question is our main interest in the current paper.
1.2 Our Contribution
In this paper, we consider the problem of finding relaxed bipartite multigraph realizations for a given degree sequence or a given partition. That is, the relaxed realizations must fulfill the degree constraints exactly but are allowed to have parallel edges. (Self-loops are disallowed.)
To evaluate the quality of a realization by a multigraph, we use two measures:
  1. (i)
    The total multiplicity of the multigraph, i.e., the number of parallel edges.
  2. (ii)
    The maximum multiplicity of the multigraph, i.e., the maximum number of edges between any two of its vertices.
As shown later, these measures are non-equivalent, in the sense that there are examples for sequences where realizations optimizing one measure are sub-optimal in the other, and vice-versa.
For relaxed realizations by general multigraphs, it follows from the characterizations given, respectively, by Owens and Trent [15] and Chungphaisan [6], how to optimize the two measurements and find the respective multigraph realizations.
For relaxed realization by bipartite multigraphs, finding a realization for BDRP (the given partition variant) that minimizes the maximum multiplicity follows from the characterization presented by Berge [14], and the BDR case (single sequence variant) was considered by us in [2]. In the current paper, we review the known literature on the problem of degree realization by general and bipartite multigraphs, strengthen the results of [2] on maximum multiplicity realizations for BDR, and also present additional results on multigraph realizations with bounded total multiplicity for BDRP.
In more detail, Section 2 introduces formally the basic notions and measures under study. We also present known results on multigraph realizations with low total multiplicity of parallel edges. The problem was solved for general graphs by Owens and Trent [15]. In Section 3 we provide a characterization for bipartite multigraphs based on a given partition (BDRP). Both characterizations are translated to Erdös-Gallai and Gale-Ryser conditions, respectively.
In Section 4 we show that optimizing total multiplicity and maximum multiplicity may lead to different realizations. Moreover, optimizing by one measure may increase the other substantially.
One necessary condition for a sequence $d=(d_{1},\ldots,d_{n})$ to be bigraphic is that it can be partitioned. If $d_{1}<n$, this problem can be decided in polynomial time. However, for a multigraph realization to exists, the inequality $d_{1}<n$ is not a necessary condition, and it follows that BDPP is NP-hard. We review this matter in greater detail in Section 5 and discuss an output sensitive algorithm to generate all partitions of a given sequence which was presented in [2]. In case the number of partitions of a sequence is small, the algorithm allows us to find optimal realizations with respect to both criteria.
In Section 6, we discuss sufficient conditions for the existence of approximate bipartite realizations that depend only on the largest degree of the sequence of a given sequence.
2 Preliminaries
Let $d=(d_{1},d_{2},\ldots,d_{n})$ be a sequence of positive integers in non-increasing order. (All sequence that we consider are assumed to be of positive integers and in a non-increasing order.) The volume of $d$ is ${\sum{d}}=\sum_{i=1}^{n}d_{i}$. For a graph $G$, denote the sequence of its vertex-degrees by $\deg(G)$. Sequence $d$ is graphic if there is a graph $G$ such that $\deg(G)=d$. We say that $G$ is a realization of $d$. Note that every realization of $d$ has $m={\sum{d}}/2$ edges. Consequently, a graphic sequence must have even volume. In turn, we call a sequence of positive integers with even volume a degree sequence. We use the operator $\circ$ to define $d\circ d^{\prime}$ as the concatenation of two degree sequences $d$ and $d^{\prime}$ (in non-increasing order).
2.1 Multigraphs as Approximate Realizations
Let $H=(V,E)$ be a multigraph without loops. In this case, $E$ is a multiset. Denote by $E_{H}(v,u)$ the multiset of edges connecting $v,u\in V$. If $\left|E_{H}(v,u)\right|>1$, we say that the edge $(v,u)$ has $\left|E_{H}(v,u)\right|-1$ excess copies. Let $E^{\prime}$ be the set that is obtained by removing excess edges from $E$. The graph $G=(V,E^{\prime})$ is called the underlying graph of $H$.
We view multigraphs as approximate realizations of sequences that are not graphic. Owens and Trent [15] gave a condition for the existence of a multigraph realization.
Theorem 1 (Owens and Trent [15]).
A degree sequence $d$ can be realized by a multigraph if and only if $d_{1}\leq\sum_{i=2}^{n}d_{i}$.
To measure the quality of an approximate realization we introduce two metrics. First, the maximum multiplicity of a multigraph $H$ is the maximum number of copies of an edge, namely
$\mbox{\sf MaxMult}(H)\triangleq\max_{(v,w)\in E}(\left|E_{H}(v,w)\right|)~{},$
and for a sequence $d$ define
$\mbox{\sf MaxMult}(d)\triangleq\min\{\mbox{\sf MaxMult}(H):H\text{ realizes }d\}~{}.$
We say that a sequence $d$ is $r$-max-graphic if $\mbox{\sf MaxMult}(d)\leq r$, for a positive integer $r$.
Second, the total multiplicity of a multigraph $H$ is the total number of excess copies, namely
$\mbox{\sf TotMult}(H)\triangleq\sum_{(v,w)\in E}(\left|E_{H}(v,w)\right|-1)=\left|E\right|-\left|E^{\prime}\right|~{},$
where $E^{\prime}$ is the edge set of the underlying graph of $H$. For a sequence $d$ define
$\mbox{\sf TotMult}(d)\triangleq\min\{\mbox{\sf TotMult}(H):H\text{ realizes }d\}~{}.$
We say a sequence $d$ is $t$-tot-graphic if $\mbox{\sf TotMult}(d)\leq t$, for a positive integer $t$.
2.2 General Multigraphs
Given a degree sequence $d$, our goal is to compute $\mbox{\sf MaxMult}(d)$ and $\mbox{\sf TotMult}(d)$.
We note that the best realization in terms of maximum multiplicity is not necessarily the same as the best one in terms of total multiplicity. See more on this issue in Section 4.
Next, we iterate the characterization of Erdös and Gallai [8] for graphic sequence.
Theorem 2 (Erdös-Gallai [8]).
A degree sequence $d$ is graphic if and only if, for $\ell=1,\ldots,n$,
$\sum_{i=1}^{\ell}d_{i}\leq\ell(\ell-1)+\sum_{i=\ell+1}^{n}\min\{\ell,d_{i}\}~{}.$ (1)
Theorem 2 implies an $\mathcal{O}(n)$ algorithm to verify whether a sequence is graphic. Chungphaisan [6] extended the above characterization to multigraphs with bounded maximum multiplicity as follows.
Theorem 3 (Chungphaisan [6]).
Let $r$ be a positive integer. A degree sequence $d$ is $r$-max-graphic if and only if, for $\ell=1,\ldots,n$,
$\sum_{i=1}^{\ell}d_{i}\leq r\ell(\ell-1)+\sum_{i=\ell+1}^{n}\min\{r\ell,d_{i}\}~{}.$ (2)
Notice the similarity to the Erdös-Gallai equations. Moreover, verify that $r\leq d_{1}$, for any $r$-max-graphic sequence $d$. It follows that $\mbox{\sf MaxMult}(d)$ can be computed in $\mathcal{O}(n\cdot\log(d_{1}))$.
The problem of finding a multigraph realization with low total multiplicity was solved by Owens and Trent [15]. They showed that the minimum total multiplicity is equal to the minimum number of degree $2$ vertices that should be added to make the sequence graphic. We provide a simpler proof of their result.
Theorem 4 (Owens and Trent [15]).
Let $d$ be a degree sequence such that $d_{1}\leq\sum_{i=2}^{n}d_{i}$. Then, $d$ is $t$-tot-graphic if and only if $d\circ 2^{t}$ is graphic.
Proof.
Let $d$ be a degree sequence such that $d_{1}\leq\sum_{i=2}^{n}d_{i}$. First, assume that $d$ can be realized by a multigraph $H$ with $\mbox{\sf TotMult}(H)\leq t$. Let $F$ be the set of excess edges in $H$. Construct a simple graph $G$ by replacing each edge $f=(x,y)\in F$ with two edges $(x,v_{f})$ and $(y,v_{f})$, where $v_{f}$ is a new vertex. Clearly, this does not change the degrees of $x$ and $y$ and adds a vertex $v_{f}$ of degree $2$. Hence the degree sequence of $G$ is $d\circ 2^{|F|}$. Also, $G$ is simple. If $|F|<t$, then one may replace any edge in $G$ with a path containing $t-|F|$ edges, yielding a graph with degree sequence $d\circ 2^{t}$.
Conversely, suppose the sequence $d\circ 2^{t}$ is graphic. Let $G$ be a simple graph that realizes the sequence. Pick a degree $2$ vertex $v$ with neighbors $x$ and $y$, replace the edges $(v,x)$ and $(v,y)$ with the edge $(x,y)$, and remove $v$ from $G$. This transformation eliminates one degree 2 vertex from $G$ without changing the remaining degrees. But it may increase the number of excess edges by one (if the edge $(x,y)$ already exists in $G$). Performing this operation for $t$ times, we obtain a multigraph $H$ with $\mbox{\sf TotMult}(H)\leq t$ and degree sequence $d$. ∎
The next corollary follows readily with Theorems 2 and 4.
Corollary 5.
Let $t$ be a positive integer, and let $d^{\prime}=d\circ 2^{t}$. Sequence $d$ is $t$-tot-graphic if and only if, for $\ell=1,\ldots,n+t$,
$\sum_{i=1}^{\ell}d^{\prime}_{i}\leq\ell(\ell-1)+\sum_{i=\ell+1}^{n+t}\min\{\ell,d^{\prime}_{i}\}~{}.$ (3)
Owens and Trent [15] implicitly suggest to compute $\mbox{\sf TotMult}(d)$ by computing the minimum $t$ such that $d\circ 2^{t}$ is graphic. Using binary search would lead to a running time of $\mathcal{O}(n\cdot\log(\mbox{\sf TotMult}(d)))$.
Several authors [18, 20] noticed that the equations of Theorem 2 are not minimal. For a degree sequence $d$ where $d_{1}>1$, let $\text{box}(d)=max\{i~{}|~{}d_{i}>i\}$. If Equation (1) holds for the index $\ell=\text{box}(d)$, then it holds for index $\ell+1$. To see this, consider the equations for the two indices and compare the change in the left hand side (LHS) and right hand side (RHS). Observe that the RHS increases at least by $(\ell+1)\cdot\ell-\ell\cdot(\ell-1)=2\ell$, while the LHS only increases by $d_{\ell+1}\leq\ell$. It follows that Equation (1) does not have to be checked for indices $\ell>\text{box}(d)$. If $d_{1}=1$, we define $\text{box}(d)=0$. Note that in this case $d$ is realized by a matching graph.
Observation 6 ([18, 20]).
A degree sequence $d$ is graphic if and only if, for $\ell=1,\ldots,\text{box}(d)$,
$\sum_{i=1}^{\ell}d_{i}\leq\ell(\ell-1)+\sum_{i=\ell+1}^{n}\min\{\ell,d_{i}\}~{}.$ (4)
On a side note, it is also known that only up to $k$ many equations have to be checked where $k$ is the number of different degrees of a sequence (cf. [14, 18, 20]).
Observation 6 helps to simplify Corollary 5.
Corollary 7.
Let $t$ be a positive integer. Degree sequence $d$ is $t$-tot-graphic if and only if, for $\ell=1,\ldots,\text{box}(d)$,
$\sum_{i=1}^{\ell}d_{i}\leq\ell(\ell-1)+\sum_{i=\ell+1}^{n}\min\{\ell,d_{i}\}+t\cdot\min\left\{\ell,2\right\}.$ (5)
Proof.
Let $d$ and $t$ be as in the corollary. In case $d_{1}=1$, the sequence $d$ is graphic, i.e., it is $t$-tot-graphic for any positive integer $t$.
Hence, assume that $d_{1}>1$. Also, let $d^{\prime}=d\circ 2^{t}$. One can verify that Equations (5) are the (reduced) Erdös-Gallai inequalities of Observation 6 for $d^{\prime}$. Moreover, $\text{box}(d)=\text{box}(d^{\prime})$, and the claim follows. ∎
Corollary 7 implies a simple algorithm to compute $\mbox{\sf TotMult}(d)$. Let
$\Delta_{\ell}(d)=\sum_{i=1}^{\ell}d_{i}-(\ell(\ell-1)+\sum_{i=\ell+1}^{n}\min\{\ell,d_{i}\}),$
for $\ell=1,\ldots,n$, be the Erdös-Gallai differences of a degree sequence $d$. Also, let $\Delta_{\text{max}}(d)=\max_{2\leq\ell\leq\text{box}(d)}\Delta_{\ell}(d)$. It follows that $t=\max\{\Delta_{1},\Delta_{\text{max}}/2\}$ implying a $\mathcal{O}(n)$ algorithm to calculate $\mbox{\sf TotMult}(d)$.
2.3 Bipartite Multigraphs
In this section, we start investigating whether a degree sequence has a bipartite realization, i.e., if it is bigraphic or not. Particularly, we are interested in multigraph realizations where the underlying graph is bipartite.
Let $d$ be a degree sequence such that ${\sum{d}}=2m$ for some integer $m$. A block of $d$ is a subsequence $a$ such that ${\sum{a}}=m$. Define the set of blocks as $B(d)$. For each $a\in B(d)$ there is a disjoint $b\in B(d)$ such that $d=a\circ b$. We call such a pair $a,b\in B(d)$ a balanced partition of $d$ since ${\sum{a}}={\sum{b}}$. Denote the set of all partitions of $d$ by ${\tt BP}(d)=\left\{\left\{a,b\right\}~{}|~{}a,b\in B(d),~{}a\circ b=d\right\}$. We say a partition $(a,b)\in{\tt BP}(d)$ is bigraphic if there is a bipartite realization $G=(A,B,E)$ of $d$ such that $\deg(A)=a$ and $\deg(B)=b$ are the vertex-degree sequences of $A$ and $B$, respectively.
Observe that, as in the case of general graphs, the best realization in terms of maximum multiplicity is not necessarily the same as the best one in terms of total multiplicity. See Section 4.
Note that not every graphic sequence has a balanced partition. Yet, if $d$ is bigraphic, then ${\tt BP}(d)$ is not empty. The Gale-Ryser theorem characterizes when a partition is bigraphic.
Theorem 8 (Gale-Ryser [9, 17]).
Let $d$ be a degree sequence and partition $(a,b)\in{\tt BP}(d)$ where $a=(a_{1},a_{2},\ldots,a_{p})$ and $b=(b_{1},b_{2},\ldots,b_{q})$. The partition $(a,b)$ is bigraphic if and only if, for $\ell=1,\ldots,p$,
$\sum_{i=1}^{\ell}a_{i}\leq\sum_{i=1}^{q}\min\{\ell,b_{i}\}~{}.$ (6)
We point out that Theorem 8 does not characterize bigraphic degree sequences. Indeed, if the partition is not specified, it is not known how to determine whether a graphic sequence is bigraphic or not. There are sequences where some partitions are bigraphic while others are not. Moreover, $|{\tt BP}(d)|$ might be exponentially large in the input size $n$.
We turn back to approximate realizations by bipartite multigraphs. A multigraph is bipartite if its underlying graph is bipartite. Analogue to above, we use the maximum and total multiplicity to measure the quality of a realization. Naturally, let
$\mbox{\sf MaxMult}^{bi}(d)\triangleq\min\{\mbox{\sf MaxMult}(H):H\text{ is bipartite and realizes }d\}~{}.$
For a partition $(a,b)\in{\tt BP}(d)$, we define
$\mbox{\sf MaxMult}^{bi}(a,b)\triangleq\min\{\mbox{\sf MaxMult}(H):H=(A,B,E)\text{ s.t. }\deg(A)=a\text{ and }\deg(B)=b\}~{}.$
Let $r$ be a positive integer. If there is a bipartite multigraph $H=(A,B,E)$ where $\mbox{\sf MaxMult}(H)\leq r$, we say that $d$ is $r$-max-bigraphic. Moreover, we say that the partition $(a,b)\in{\tt BP}(d)$, where $a=\deg(A)$ and $b=\deg(B)$, is $r$-max-bigraphic. Miller [14] cites the following result of Berge characterizing $r$-max-bigraphic partitions.
Theorem 9 (Berge [14]).
Let $r$ be a positive integer. Consider a degree sequence $d$ and a partition $(a,b)\in{\tt BP}(d)$, where $a=(a_{1},\ldots,a_{p})$ and $b=(b_{1},\ldots,b_{q})$. Then $(a,b)$ is $r$-max-bigraphic if and only if, for $\ell=1,\ldots,p$,
$\sum_{i=1}^{\ell}a_{i}\leq\sum_{i=1}^{q}\min\{\ell r,b_{i}\}~{}.$ (7)
Note the similarity to the Gale-Ryser theorem. Theorem 9 implies that $\mbox{\sf MaxMult}^{bi}(a,b)$ can be computed in $\mathcal{O}(n\cdot\log(d_{1}))$ using binary search.
For the second approximation criteria, we bound the total multiplicity of a bipartite multigraph realization. Define
$\mbox{\sf TotMult}^{bi}(d)\triangleq\min\{\mbox{\sf TotMult}(H):H\text{ is bipartite and realizes }d\}~{}.$
Additionally, for a partition $(a,b)\in{\tt BP}(d)$, we define
$\mbox{\sf TotMult}^{bi}(a,b)\triangleq\min\{\mbox{\sf TotMult}(H):H=(A,B,E)\text{ s.t. }\deg(A)=a\text{ and }\deg(B)=b\}~{}.$
We present our results on determining $\mbox{\sf TotMult}^{bi}(a,b)$ in the next section. In Sections 5 and 6, we consider $\mbox{\sf MaxMult}^{bi}(d)$ and $\mbox{\sf TotMult}^{bi}(d)$.
3 Multigraph Realizations of Bi-sequences
In this section, we are interested in bipartite multigraph realizations with low total multiplicity, assuming that we are given a sequence and a specific balanced partition. First, we provide a characterization similar to Theorem 4 for bipartite multigraph realizations for a given partition.
Theorem 10.
Let $d$ be a degree sequence and $t$ be a positive integer. Then, $d$ is $t$-tot-bigraphic if and only if there exists a partition $(a,b)\in{\tt BP}(d)$ such that $(a\circ 1^{t},b\circ 1^{t})$ is bigraphic.
Proof.
Let $d$, $t$ be as in the theorem. Assume that there is a bipartite multigraph $H=(L,R,E)$ with $\mbox{\sf TotMult}(H)\leq t$. Hence, there is a partition $(a,b)\in{\tt BP}(d)$ where $\deg(L)=a$ and $\deg(R)=b$. Let $F$ be the set of excess edges in $H$. Construct a bipartite graph $G$ by applying the following transformation. For every excess edge $(x,y)\in F$, add a new vertex $x_{e}$ to $A$ and a new vertex $y_{e}$ to $B$, and replace $(x,y)$ by the two edges $(x,y_{e})$ and $(x_{e},y)$. Note that $x_{e}$ and $y_{e}$ are placed on opposite partitions of $G$. Since there are $t$ excess edges, $G$ realizes $(a\circ 1^{t},b\circ 1^{t})$ without excess edges.
For the other direction, assume that there exists a partition $(a,b)\in{\tt BP}(d)$ such that $(a\circ 1^{t},b\circ 1^{t})$ is realized by a bipartite graph $G=(L,R,E)$. Let $x_{1},\ldots,x_{t}$ and $y_{1},\ldots,y_{t}$ be some vertices of degree one in $L$ and $R$, respectively. Also, for every $i$, let $y^{\prime}_{i}$ (respectively, $x^{\prime}_{i}$) be the only neighbor of $x_{i}$ (resp., $y_{i}$). Construct a bipartite multigraph $H$ by replacing the edges $(x_{i},y^{\prime}_{i})$ and $(x^{\prime}_{i},y_{i})$ with the edge $(x^{\prime}_{i},y^{\prime}_{i})$ and discarding the vertices $x_{i}$ and $y_{i}$, for every $i$. Since this transformation may add up to $t$ excess edges, we have that $\mbox{\sf TotMult}(H)\leq t$. ∎
The above characterization leads to extended Gale-Ryser conditions for total multiplicity.
Theorem 11.
Let $d$ be a degree sequence with partition $(a,b)\in{\tt BP}(d)$ where $a=(a_{1},\ldots,a_{p})$ and $b=(b_{1},\ldots,b_{q})$, and let $t$ be a positive integer. The partition $(a,b)$ is $t$-tot-bigraphic if and only if, for all $\ell\in 1,\ldots,p$,
$\sum_{i=1}^{\ell}a_{i}\leq\sum_{i=1}^{q}\min\{\ell,b_{i}\}+t~{}.$ (8)
Proof.
Consider $(a,b)$ and $t$ as in the theorem. One can verify that the following equations are the Gale-Ryser conditions of Theorem 8 for the partition $(a,b)$: For all $\ell\in 1,\ldots,p$,
$\sum_{i=1}^{\ell}a_{i}\leq\sum_{i=1}^{q}\min\{\ell,b_{i}\}+t~{},$ (9)
and for all $h\in 1,\ldots,t$,
$\sum_{i=1}^{p}a_{i}+h\leq\sum_{i=1}^{q}\min\{p+h,b_{i}\}+t~{}.$ (10)
To finish the proof, we argue that Equation 10 holds for any $h\in\left\{0,\ldots,t\right\}$ if Equation 9 holds for $\ell=p$. Recall that $\sum_{i=1}^{q}\min\{p+h,b_{i}\}=\sum_{i=1}^{q}b_{i}$ if $p+h\geq b_{1}$. It follows that Equation (10) holds for indices $h\geq b_{1}-p$.
Observe that $\sum_{i=1}^{q}\min\{p+h+1,b_{i}\}-\sum_{i=1}^{q}\min\{p+h,b_{i}\}\geq 1$ for $p+h<b_{1}$, i.e., the RHS of Equation (10) grows by at least $1$ when moving from index $p+h$ to index $p+h+1$. By assumption, Equation (9) holds for $\ell=p$, implying that Equation (10) holds for $h=0$. Since the LHS of Equation (10) grows by $1$ exactly, Equation (10) holds for indices $h<b_{1}-p$. ∎
Given a degree sequence $d$ with partition $(a,b)\in{\tt BP}(d)$, Theorem 11 implies that
$\mbox{\sf TotMult}^{bi}(a,b)=\max_{1\leq\ell\leq p}\left(\sum_{i=1}^{\ell}a_{i}-\sum_{i=1}^{q}\min\{\ell,b_{i}\}\right)~{}.$
It follows that $\mbox{\sf TotMult}^{bi}(a,b)$ can be computed in time $\mathcal{O}(n)$.
4 Total Multiplicity vs. Maximum Multiplicity
In this section we show that the measures of total multiplicity and maximum multiplicity sometimes display radically different behavior. Specifically, we show that there are sequences such that a realization that minimizes the total multiplicity may be far from achieving minimum maximum multiplicity, and vice versa.
First, we notice that by definition of TotMult and $\mbox{\sf TotMult}^{bi}$, in order to minimize the total multiplicity one needs to use a maximum number of edges, or to maximize the number of edges in the underlying graph.
Observation 12.
Let $d$ be a sequence and let $H=(V,E)$ be a multigraph that realizes $d$. Let $G^{\prime}=(V,E^{\prime})$ be the underlying graph of $H$. Then,
  • $\mbox{\sf TotMult}(H)=\mbox{\sf TotMult}(d)$ if and only if $|E^{\prime}|$ is maximized.
  • $\mbox{\sf TotMult}^{bi}(H)=\mbox{\sf TotMult}^{bi}(d)$ if and only if $|E^{\prime}|$ is maximized.
4.1 General Graphs
Let $n\geq 5$, and consider the sequence
${\hat{d}}=\left((2n-2)^{2},(n-1)^{n-2}\right)~{}.$
Observe that $\sum{\hat{d}}=(n-1)(n+2)$. The following two lemmas show that a realization of ${\hat{d}}$ attaining minimum total multiplicity is far from obtaining minimum maximum multiplicity, and vise versa.
Lemma 13.
$\mbox{\sf TotMult}({\hat{d}})=n-1$, and if $H$ realizes ${\hat{d}}$ such that $\mbox{\sf TotMult}(H)=\mbox{\sf TotMult}({\hat{d}})$, then $\mbox{\sf MaxMult}(H)=n$.
Proof.
Consider a mutligraph $H$ which is composed of a full graph and $n-1$ copies of the edge $(1,2)$. More formally, let $H=(V,E)$ be a multigraph where
$E=\left\{(i,j)\mid 1\leq i<j\leq n\right\}\uplus\biguplus_{t=1}^{n-1}\left\{(1,2)\right\}~{}.$
Recall that $E$ is a multiset. See example in Figure 1a.
We have that $\deg(1)=\deg(2)=(n-1)+(n-1)=2n-2$, and $\deg(i)=(n-1)$, for $i>2$, as required. Thus, $H$ realizes ${\hat{d}}$. Observe that $\left|E^{\prime}\right|=\binom{n}{2}$, hence by 12 we have that
$\mbox{\sf TotMult}({\hat{d}})=\mbox{\sf TotMult}(H)=(n-1)~{}.$
Let $\bar{H}$ be a realization such that $\mbox{\sf TotMult}(\bar{H})=\mbox{\sf TotMult}({\hat{d}})$. It follows that $\left|\bar{E}^{\prime}\right|=\binom{n}{2}$. Consider a vertex $i$, such that $i>2$. All edges touching $i$ must be used at least once. Since ${\hat{d}}_{i}=n-1$, all edge adjacent to $i$ must be used exactly once. It follows that all excess edges are connected to the vertices $\left\{1,2\right\}$. It follows that $\bar{H}=H$. Hence, $H$ minimizes the maximum multiplicity, and $\mbox{\sf MaxMult}(H)=n$. ∎
(a) Optimal TotMult realization $H_{1}$.
(b) Optimal MaxMult realization $H_{2}$.
Figure 1: Optimal multigraph realizations for the sequence ${\hat{d}}=(8^{2},4^{3})$ ($n=5$). On the left we have $\mbox{\sf TotMult}(H_{1})=4$ and $\mbox{\sf MaxMult}(H_{1})=5$, while on the right we have $\mbox{\sf TotMult}(H_{2})=7$ and $\mbox{\sf MaxMult}(H_{2})=2$.
Lemma 14.
$\mbox{\sf MaxMult}({\hat{d}})=2$, and if $H$ realizes ${\hat{d}}$ such that $\mbox{\sf MaxMult}(H)=\mbox{\sf MaxMult}({\hat{d}})$, then $\mbox{\sf TotMult}(H)\geq 2n-3$.
Proof.
Consider a vertex $1$ (or $2$). To minimize its load, its degree requirement should be distributed equally among the rest of the vertices. This leads to a realization $H$ in which each edge of $1$ and $2$ has two copies. The degree requirements of the rest of the vertices are obtained by removing a cycle from a complete graphs (this is the reason for requiring $n\geq 5$). Formally, $H=(V,E)$ is define as follows:
$E=\left\{(1,i),(1,i)\mid i\geq 2\right\}\uplus\left\{(2,i),(2,i)\mid i\geq 3\right\}\uplus\left\{(i,j)\mid 2<i,j\neq i+1\text{ and }(i,j)\neq(3,n)\right\}~{}.$
See example in Figure 1b.
We have that $\deg(1)=\deg(2)=2(n-1)$, and $\deg(i)=2+(n-1-2)=n-1$, for $i>2$, as required. Thus, $H$ realizes ${\hat{d}}$. Moreover, each edges has at most two copies, which means that $\mbox{\sf MaxMult}({\hat{d}})=2$.
Observe that an edge $(i,j)$, where $i,j>2$ has at most a single copy. Hence, $H$ minimizes the total multiplicity. In addition, $\mbox{\sf TotMult}(H)=(n-1)+(n-2)=2n-3$. ∎
Corollary 15.
Let $n\geq 5$. There exists a sequence ${\hat{d}}$ of length $n$ such that $\mbox{\sf TotMult}(H_{2})-\mbox{\sf TotMult}(H_{1})=n-2$ and $\mbox{\sf MaxMult}(H_{1})-\mbox{\sf MaxMult}(H_{2})=n-2$, for any $H_{1}$ and $H_{2}$ such that $\mbox{\sf TotMult}(H_{1})=\mbox{\sf TotMult}({\hat{d}})$ and $\mbox{\sf MaxMult}(H_{2})=\mbox{\sf MaxMult}({\hat{d}})$.
4.2 Bipartite Graphs
Let $n$ be an even integer such that $n\geq 4$, and consider the sequence
${\tilde{d}}=\left(n^{2},(n/2)^{n-2}\right)~{}.$
Lemma 16.
$\mbox{\sf TotMult}^{bi}({\tilde{d}})=n/2$, and if $H$ realizes ${\tilde{d}}$ such that $\mbox{\sf TotMult}^{bi}(H)=\mbox{\sf TotMult}^{bi}({\tilde{d}})$, then $\mbox{\sf MaxMult}^{bi}(H)\geq n/2+1$.
Proof.
Let $a=b=(n,(\frac{n}{2})^{(n-2)/2})$. We construct a multigraph $H$ that realizes $(a,b)$, which consist of a complete bipartite graph and $n/2$ copies of the edge $(1,1)$. Formally, $H=(A,B,E)$, where
$E=\left\{(i,j)\mid 1\leq i,j\leq n/2\right\}\uplus\biguplus_{t=1}^{n/2}\left\{(1,1)\right\}~{}.$
See example in Figure 2a.
On both sides we have that $\deg(1)=\frac{n}{2}+\frac{n}{2}=n$, and $\deg(i)=\frac{n}{2}$, for $i>1$, as required. Thus, $H$ realizes $(a,b)$. Observe that $\left|E^{\prime}\right|=\frac{n}{2}\cdot\frac{n}{2}$, hence by 12 we have that
$\mbox{\sf TotMult}^{bi}({\tilde{d}})=\mbox{\sf TotMult}^{bi}(a,b)=\mbox{\sf TotMult}^{bi}(H)=n/2~{}.$
Let $\bar{H}$ be a realization such that $\mbox{\sf TotMult}^{bi}(\bar{H})=\mbox{\sf TotMult}^{bi}(a,b)$. It follows that $\left|\bar{E}^{\prime}\right|=\binom{n}{2}$. Consider a vertex $i$, such that $i>2$. All edges touching $i$ must be used at least once. Since ${\tilde{d}}_{i}=\frac{n}{2}$, all edges touching $i$ must be used exactly once. It follows that all excess edges are connected to the vertices $\left\{1,2\right\}$. Hence, $\bar{H}=H$. Also, $\mbox{\sf MaxMult}^{bi}(H)=n/2+1$. ∎
(a) Optimal $\mbox{\sf TotMult}^{bi}$ realization $H_{1}$.
(b) Optimal $\mbox{\sf MaxMult}^{bi}$ realization $H_{2}$.
(c) Optimal $\mbox{\sf MaxMult}^{bi}$ realization $H_{3}$.
Figure 2: Multigraph bipartite realizations for the sequence $d=(6^{2},3^{4})$. On the left we have $\mbox{\sf TotMult}^{bi}(H_{1})=3$ and $\mbox{\sf MaxMult}^{bi}(H_{1})=4$; In the center we have $\mbox{\sf TotMult}^{bi}(H_{2})=5$ and $\mbox{\sf MaxMult}^{bi}(H_{1})=2$; On the right we have $\mbox{\sf TotMult}^{bi}(H_{2})=4$ and $\mbox{\sf MaxMult}^{bi}(H_{2})=2$.
Lemma 17.
$\mbox{\sf MaxMult}^{bi}({\tilde{d}})=2$, and if $H$ realizes ${\tilde{d}}$ such that $\mbox{\sf MaxMult}(H)=\mbox{\sf MaxMult}({\tilde{d}})$, then $\mbox{\sf TotMult}(H)\geq n-2$.
Proof.
There are two possible partitions for ${\tilde{d}}$:
(P1) $a=b=\left(n,(\frac{n}{2})^{(n-2)/2}\right)$, and
(P2) $a^{\prime}=\left(n^{2},(\frac{n}{2})^{n/2-3}\right)$ and $b^{\prime}=\left((\frac{n}{2})^{n/2+1}\right)$.
We first consider partition (P1). Consider a vertex $1\in A$ (or $1\in B$). To minimize its load, its degree requirement should be distributed equally among the rest of the vertices on the other side of the partition. This leads to a realization $H$, where vertex $1\in A$ is connected to all the vertices in $B$ by two copies, while vertex $1\in B$ is connected to all the vertices in $A$ by two copies. The requirement of the other vertices is obtained using a complete bipartite graph minus a perfect matching. Hence, $H=(A,B,W)$, where
$E=\left\{(1,1),(1,1)\right\}\uplus\left\{(1,i),(1,i),(i,1),(i,1)\mid i\geq 2\right\}\uplus\left\{(i,j)\mid i,j\geq 2,j\neq i\right\}~{}.$
See example in Figure 2b.
We have that $\deg(1)=2(n/2)=n$, and $\deg(i)=2+(n/2-2)=n/2$, for $i>2$, on both sides, as required. Thus, $H$ realizes $(a,b)$. Moreover, $\mbox{\sf MaxMult}^{bi}(a,b)=\mbox{\sf MaxMult}^{bi}(H)=2$, since each edge has at most two copies. Observe that an edge $(i,j)$, where $i,j>1$ has at most a single copy. Hence, $H$ minimizes the maximum multiplicity. In addition, $\mbox{\sf TotMult}(H)=n/2+n/2-1=n-1$.
Next, consider partition (P2). We construct a realization $H^{\prime}$ as follows:
$E=\left\{(i,j)\mid i\leq 2\text{ or }i>2\text{ and }j\neq i\right\}\uplus\left\{(1,j)\mid j\leq n/2-1\right\}\uplus\left\{(2,j)\mid j\geq 3\right\}~{}.$
See example in Figure 2c.
On the left side, we have that $\deg(1)=\deg(2)=n/2+1+n/2-1=n$, and $\deg(i)=n/2+1-1=n/2$, for $i>3$. On the right side, $\deg(j)=n/2-1+2$, for $i\in\left\{3,\ldots,n/2-1\right\}$, and $\deg(j)=n/2+1$, for $i\in\left\{1,2,n/2,n/2+1\right\}$. Thus, $H^{\prime}$ realizes $(a^{\prime},b^{\prime})$. Moreover, $\mbox{\sf MaxMult}^{bi}(a^{\prime},b^{\prime})=\mbox{\sf MaxMult}^{bi}(H^{\prime})=2$, since each edge has at most two copies. Furthermore, $\mbox{\sf TotMult}(H^{\prime})=2(n/2-1)=n-2$. ∎
Corollary 18.
Let $n$ be an even integer such that $n\geq 4$. There exists a sequence ${\tilde{d}}$ of length $n$ such that $\mbox{\sf TotMult}^{bi}(H_{2})-\mbox{\sf TotMult}^{bi}(H_{1})=n/2-2$ and $\mbox{\sf MaxMult}^{bi}(H_{1})-\mbox{\sf MaxMult}^{bi}(H_{2})=n/2-1$, for any $H_{1}$ and $H_{2}$ such that $\mbox{\sf TotMult}^{bi}(H_{1})=\mbox{\sf TotMult}^{bi}({\tilde{d}})$ and $\mbox{\sf MaxMult}^{bi}(H_{2})=\mbox{\sf MaxMult}^{bi}({\tilde{d}})$.
5 Bipartite Realization of a Single Sequence
In this section, we study the following question: given a degree sequence $d$, can it be realized as a multigraph whose underlying graph is bipartite? Also, if there exists such a realization, we would like to find one which minimizes the maximum or the total multiplicity.
5.1 Hardness Result
Given a sequence and a balanced partition one may construct a bipartite multigraph realization by assigning edges in an arbitrary manner.
Observation 19.
Let $d$ be a sequence and let $(\ell,r)\in{\tt BP}(d)$ be a partition of $d$. Then, there exists a bipartite multigraph realization of $(\ell,r)$.
It follows that deciding whether a degree sequence $d$ can it be realized as a multigraph whose underlying graph is bipartite is NP-hard.
Theorem 20.
Deciding if a degree sequence $d$ admits a bipartite multigraph realization is NP-hard.
Proof.
We prove the theorem using a reduction from the Partition problem. Recall that Partition contains all sequences $(a_{1},\ldots,a_{n})$ such that there exists an index set $S\subseteq[1,n]$ for which $\sum_{i\in S}a_{i}=\sum_{i\not\in S}a_{i}$ (see. e.g., [10]). Observation 19 implies that $d$ is a Partition instance if and only if $d$ admits a bipartite mulitgraph realization. ∎
Since Partition admits a pseudo-polynomial time algorithm, we have the following.
Theorem 21.
Deciding if a sequence $d$ admits a bipartite multigraph realization can be done in pseudo-polynomial time.
5.2 Computing all Balanced Partitions of a Degree Sequence
We describe an algorithm that given a degree sequence $d$, computes all balanced partitions of $d$. The algorithm relies on the self-reducability of the Subset-Sums problem. Recall that in Subset-Sum the input is a sequence of numbers $(a_{1},\ldots,a_{n})$ and an additional number $t$, and the question is whether there is a subset $S$ such that $\sum_{i\in S}a_{i}=t$. Let Subset-Sums-DP be a dynamic programming algorithm for Subset-Sum whose running time is denoted by $T_{DP}(a,t)$ (see, e.g., [7]). The running time of the dynamic programming algorithm can be bounded by $O(n\cdot\min\left\{\sum a,\left|{\tt BP}(a)\right|\right\})$.
In this section we abuse notation by presenting a sequence $d$ as a sequence of $q$ blocks, namely $d=(d_{1}^{n_{1}},d_{2}^{n_{2}},\ldots,d_{q}^{n_{q}})$. The algorithm for computing all balanced partitions of a sequence $d$ is recursive, and it works as follows. The input is a suffix of $d$, i.e., $(d_{k}^{n_{k}},\ldots,d_{q}^{n_{q}})$, represented by $d$ and $k$, and a partition $(L,R)$ of the prefix $(d_{1}^{n_{1}},\ldots,d_{k-1}^{n_{k-1}})$. If the current suffix is empty, then it checks whether the current partition is balanced, and if it is balanced, then the partitioned is returned. Otherwise, it checks whether the current partition can be completed to a balanced partition. If the answer is YES, then the algorithm is invoked for the $n_{k}+1$ options of adding the $n_{k}$ copies of $d_{k}$ to $(L,R)$. The initial call is $(d,1,\emptyset,\emptyset)$.
1 if $k=q+1$ then
2       if $(L,R)\in{\tt BP}(d)$ then
3             return $\left\{(L,R)\right\}$
4
[…]
6       return $\emptyset$
7 $\mathcal{P}\leftarrow\emptyset$
8 for $i=0$ to $n_{k}$ do
9       $L^{\prime}\leftarrow L\circ(d_{k}^{i})$
10       $R^{\prime}\leftarrow L\circ(d_{k}^{n_{k}-i})$
11       $\mathcal{P}\leftarrow\mathcal{P}\cup\text{Partitions}(d,k+1,L^{\prime},R^{\prime})$
return $\mathcal{P}$
Algorithm 1 Partitions$(d,k,L,R)$
Lemma 22.
Algorithm Partitions returns all balanced partitions of $d$.
Proof.
Observe that each recursive call of the algorithm corresponds to a partition of a prefix of $d$. We prove that, given a prefix partition, the algorithm returns all of its balanced completions.
At the recursion base, if $(L,R)$ is a partition of $d$, then it is returned if and only if $(L,R)\in{\tt BP}(d)$. For the inductive step, let $(L,R)$ be a partition of the prefix $(d_{1}^{n_{1}},\ldots,d_{k-1}^{n_{k-1}})$. If $(L,R)$ gets a NO from Subset-Sums-Alg, then it cannot be completed to a balanced partition, and indeed no partition that corresponds to the prefix $(L,R)$ is returned. If $(L,R)$ gets a YES, then all possible partitions of $(d_{1}^{n_{1}},\ldots,d_{k}^{n_{k}})$ are checked. By the inductive hypothesis the algorithm returns all balanced partitions that complete $(d_{1}^{n_{1}},\ldots,d_{k-1}^{n_{k-1}})$. ∎
The complexity of Algorithm Partitions is dominated by the total time spent on the invocations of Subset-Sums-DP, therefore we need to bound the number of invocations of Subset-Sums-DP. More specifically, we show the following bound.
Lemma 23.
Algorithm Partition invokes Subset-Sums-DP at most $2n\cdot\left|{\tt BP}(d)\right|$ times.
Proof.
Let us illustrate the recursive execution of the algorithm on $d$ by a computation tree $T$ consisting of $q+1$ levels. Each node in the tree is labeled by a triple $(L,R,A)$, where $A\in\{YES,NO\}$, The first two entries in the label corresponds to the prefix partition $(L,R)$ in the invocation, and the third corresponds to whether the partition can be completed to a balanced partition. Note that each such node corresponds to a single invocation of Subset-Sums-DP (or alternatively checking whether $\sum L=\sum R$ in level $q+1$). It follows that the number of invocations of Subset-Sums-DP is bounded by the size of the computation tree, not including level $q+1$.
We refer to a node as a YES-node (NO-node) if its label end with a YES (NO). Observe that all NO-nodes are leaves. On the other hand, there may be internal YES-nodes. If a YES-node is a leaf, then it corresponds to a balanced partition $(L,R)$. Clearly, the number of YES-nodes in level $k+1$ of the tree is no less than the number of YES-nodes in level $k+1$. Moreover, a NO-node must have a YES-node as a sibling, hence the number of NO-nodes in level $k\leq q$ is at most $n_{k}$ times the number of YES-nodes in level $k$. Adding it all up we get:
$\sum_{k=1}^{q}\left|{\tt BP}(d)\right|(n_{k}+1)=(n+q)\left|{\tt BP}(d)\right|\leq 2n\cdot\left|{\tt BP}(d)\right|~{}.$
Hence, the lemma allows us to get an upper bound on the time complexity of the algorithm.
Corollary 24.
The running time of Algorithm Partition is $O(n^{2}\left|{\tt BP}(d)\right|\min\left\{\sum a,\left|{\tt BP}(a)\right|\right\})$.
Clearly, the above running time becomes polynomial, if $\left|{\tt BP}(d)\right|$ is polynomial.
Due to Theorem 9, the minimum $r$ such that a given partition is $r$-max-bigraphic can be computed efficiently implying the following result.
Corollary 25.
Let $d$ be a degree sequence of length $n$ such that $\left|{\tt BP}(d)\right|=\mathcal{O}(n^{c})$, for some constant $c$. Then, $\mbox{\sf MaxMult}^{bi}(d)$ can be computed in polynomial time.
Similarly, Theorem 11 implies the following.
Corollary 26.
Let $d$ be a degree sequence of length $n$ such that $\left|{\tt BP}(d)\right|=\mathcal{O}(n^{c})$, for some constant $c$. Then, $\mbox{\sf TotMult}^{bi}(d)$ can be computed in polynomial time.
We remark that a useful special subclass consists of sequences with a constant number of different degrees, since such a sequence can have at most polynomially many different partitions.
Corollary 27.
Let $q$ be some constant and $d=(d_{1}^{n_{1}},d_{2}^{n_{2}},\ldots,d_{q}^{n_{q}})$ be a degree sequence, where $n=\sum_{i=1}^{q}n_{i}$. Then, ${\tt BP}(d)=\mathcal{O}(n^{c})$, for some constant $c$.
6 Small Maximum Degree Sequences
Towards attacking the realizability problem of general bigraphic sequences, we look at the question of bounding the total deviation of a nonincreasing sequence $d=(d_{1},\ldots,d_{n})$ as a function of its maximum degree, denoted $\Delta=d_{1}$.
Burstein and Rubin [5] consider the realization problem for directed graphs with loops, which is equivalent to BDRP. (Directed edges go from the first partition to the second.) They give the following sufficient condition for a pair of sequences to be the in- and out-degrees of a directed graph with loops.
Theorem 28 (Burstein and Rubin [5]).
Consider a degree sequence $d$ with a partition $(a,b)\in{\tt BP}(d)$ assuming that $a$ and $b$ have the same length $p$. Let $\sum a=\sum b=pc$ where $c$ is the average degree. If $a_{1}b_{1}\leq pc+1$, then $d$ is realizable by a directed graph with loops.
In what follows we make use of the following straightforward technical claim which slightly strengthens a similar claim from [2].
Observation 29.
Consider a nonincreasing integer sequence $d=(d_{1},\ldots,d_{k})$ of total sum $\sum d=D$. Then, ${\sum{({{d}[\ell]})}}\geq\left\lceil\ell D/k\right\rceil$, for every $1\leq\ell\leq k$.
Proof.
Since $d$ is nonincreasing, $\frac{1}{\ell}\sum_{i=1}^{\ell}d_{i}\geq\frac{1}{k-\ell}\sum_{i=\ell+1}^{k}d_{i}~{}.$ Consequently,
$D=\sum_{i=1}^{k}d_{i}=\sum_{i=1}^{\ell}d_{i}+\sum_{i=\ell+1}^{k}d_{i}\geq\sum_{i=1}^{\ell}d_{i}+\left\lceil\frac{k-\ell}{\ell}\sum_{i=1}^{\ell}d_{i}\right\rceil=\left\lceil\frac{k}{\ell}\sum_{i=1}^{\ell}d_{i}\right\rceil~{},$
implying the claim. ∎
6.1 Bounding the Maximum Multiplicity
Theorem 28 is extended to bipartite multigraphs with bounded maximum multiplicity, i.e., to $r$-max-bigraphic sequences. The following is a slightly stronger version of Lemma 14 from [2].
Lemma 30.
Let $r$ be a positive integer. Consider a degree sequence $d$ of length $n$ with a partition $(a,b)\in{\tt BP}(d)$. If $a_{1}\cdot b_{1}\leq r\cdot\sum d/2+r$, then $(a,b)$ is $r$-max-bigraphic.
Proof.
Let $r$, $d$ and $(a,b)$ as in the lemma where $a=(a_{1},a_{2},\ldots,a_{p})$ and $b=(b_{1},b_{2},\ldots,b_{q})$. Moreover, let $X=\sum a=\sum b=\sum d/2$. To prove the claim, we assume that $a_{1}\cdot b_{1}\leq r\cdot X+r$, and show that Equation (7) holds for a fixed index $\ell\in[p]$. The lemma then follows due to Theorem 9.
First, we consider the case where $b_{1}\leq\ell r$. Then, $\sum_{i=1}^{q}\min\{\ell r,b_{i}\}=X\geq\sum_{i=1}^{\ell}a_{i}$, and Equation (7) holds.
In the following, we assume that $\ell r<b_{1}$. Note that the conjugate sequence $\widetilde{b}$ of $b$ is nonincreasing, and that $\sum_{j=1}^{\ell r}\widetilde{b}_{j}=\sum_{i=1}^{q}\min\{\ell r,b_{i}\}$. By Observation 29,
$\sum_{i=1}^{q}\min\{\ell r,b_{i}\}\geq\left\lceil\ell rX/b_{1}\right\rceil\geq\left\lceil\ell(a_{1}b_{1}-r)/b_{1}\right\rceil=\left\lceil\ell a_{1}-\ell r/b_{1}\right\rceil=\ell a_{1}~{}.$
As $a$ is nonincreasing, we have that $\sum_{i=1}^{\ell}a_{i}\leq\ell a_{1}\leq\sum_{i=1}^{q}\min\{\ell r,b_{i}\}$. The lemma follows. ∎
Lemma 31.
There exists a degree sequence $d$ with a partition $(a,b)\in{\tt BP}(d)$, such that $a_{1}\cdot b_{1}=r\cdot\sum d/2+r$, which is $r$-max-bigraphic, but not $(r-1)$-max-bigraphic.
Proof.
Consider the sequence $d=(q^{2k-1},(q-1)^{2})$ for positive integers $q,k$ such that $q=r\cdot k$. This sequence has a unique partition $(a,b)\in{\tt BP}(d)$, where $a=b=(q^{k-1},(q-1))$. One can verify that $a_{1}b_{1}=q^{2}$, while
$r\cdot\sum d/2+r=r(qk-1)+r=rqk=q^{2}~{}.$
The partition $(a,b)$ is $r$-max-bigraphic, but no better. ∎
Lemma 30 is stated for a given partition (BDRP). For BDR, we immediately have the following which is a slight improvement over Corollary 16 form [2].
Corollary 32.
Let $r$ be a positive integer and $d$ be a partitionable degree sequence. If $d_{1}^{2}\leq r\cdot\sum d/2+r$, then any partition $(a,b)\in{\tt BP}(d)$ is $r$-max-bigraphic.
6.2 Bounding the Total Multiplicity
In this section, we establish results for total multiplicity analogous to those obtained in the previous section for the maximum multiplicity.
Lemma 33.
Let $t$ be a positive integer. Consider a degree sequence $d$ of length $n$ with a partition $(a,b)\in{\tt BP}(d)$. If $a_{1}\cdot b_{1}\leq\sum d/2+t+1$, then $(a,b)$ is $t$-tot-bigraphic.
Proof.
Let $t$, $d$ and $(a,b)$ as in the lemma where $a=(a_{1},a_{2},\ldots,a_{p})$ and $b=(b_{1},b_{2},\ldots,b_{q})$, and let $X=\sum a=\sum b=\sum d/2$. To prove the claim, we assume that $a_{1}\cdot b_{1}\leq X+t+1$, and show that Equation (8) holds for every index $\ell\in[p]$. The lemma then follows due to Theorem 11.
First, consider the case where $\ell\geq b_{1}$. In this case,
$\sum_{i=1}^{q}\min\{\ell,b_{i}\}=\sum b=X\geq\sum_{i=1}^{\ell}a_{i}~{},$
and Equation (8) holds.
Next, assume that $\ell<b_{1}$. Note that the conjugate sequence $\widetilde{b}$ of $b$ is nonincreasing, and that $\sum_{j=1}^{\ell}\widetilde{b}_{j}=\sum_{i=1}^{q}\min\{\ell,b_{i}\}$. By Observation 29,
$\sum_{i=1}^{q}\min\{\ell,b_{i}\}+t\geq\left\lceil\frac{\ell X}{b_{1}}\right\rceil+t\geq\left\lceil\frac{\ell(a_{1}b_{1}-t-1)}{b_{1}}\right\rceil+t=\left\lceil\ell a_{1}-\frac{\ell(t+1)}{b_{1}}\right\rceil+t\geq\ell a_{1}~{}.$
As $a$ is nonincreasing, we have that $\sum_{i=1}^{\ell}a_{i}\leq\ell a_{1}\leq\sum_{i=1}^{q}\min\{\ell,b_{i}\}+t$. The lemma follows. ∎
The following lemma shows that the above bound it tight.
Lemma 34.
There exists a degree sequence $d$ with a partition $(a,b)\in{\tt BP}(d)$, such that $a_{1}\cdot b_{1}=\sum d/2+t+2$, and $(a,b)$ is not $t$-tot-bigraphic.
Proof.
Consider the sequence $d=(k^{2(k-1)},1^{2})$, for a positive integer $k>1$. This sequence has only one partition $(a,b)\in{\tt BP}(d)$, where $a=b=(k^{k-1},1)$. Observe that $a_{1}b_{1}=k^{2}$, while $\sum d/2=k(k-1)+1$.
Assume that $t=k-2$. Hence, $a_{1}b_{1}=\sum d/2+t+1$. For every $\ell<k$, we have that
$\sum_{i=1}^{k}\min\{\ell,b_{i}\}+t=k+(\ell-1)(k-1)+k-2=\ell k-\ell-1+k\geq\ell k=\sum_{i=1}^{\ell}a_{i}~{}.$
For $\ell=k$, we have
$\sum_{i=1}^{k}\min\{\ell,b_{i}\}+t\geq\sum d/2=\sum_{i=1}^{k}a_{i}~{}.$
Now assume that $t=k-3$. Hence, $a_{1}b_{1}=\sum d/2+t+2$. For every $\ell<k$, we have that
$\sum_{i=1}^{k}\min\{\ell,b_{i}\}+t=k+(\ell-1)(k-1)+k-3=\ell k-\ell-2+k~{}.$
If $\ell=k-1$, we get that
$\sum_{i=1}^{k}\min\{\ell,b_{i}\}+t=(k-1)k-(k-1)-2+k=(k-1)k-1<\sum_{i=1}^{\ell}a_{i}~{},$
which means that $(a,b)$ is not $t$-tot-bigraphic. ∎
Similar to above, Lemma 34 (stated for BDRP) implies the following for BDR.
Corollary 35.
Let $t$ be a positive integer and $d$ be a partitionable degree sequence. If $d_{1}^{2}\leq\sum d/2+t+1$, then any partition $(a,b)\in{\tt BP}(d)$ is $t$-tot-bigraphic.